suppose the annual level of aggregate demand is given by AD=40+0.8Y where Y is annual GDP. Find out the amount of daily rate of involuntary change in inventory, when producers planned income level is (i) Y=380 and (ii) Y= 520

Suppose that a country devalues its currency by 13% but at the same time domestic prices rise by 11%. What real devaluation has been achieved? What if prices have risen by 16%?
Answer #1

(1)

Unplanned inventory per year = AD - Y

Unplanned inventory per day = (AD - Y) / 365, assuming 365 days a year.

AD = 40 + 0.8Y

(i) When Y = 380,

AD = 40 + (0.8 x 380) = 40 + 304 = 344

Unplanned inventory per day = (344 - 380) / 365 = -36/365 = -0.0986 (a negative value means inventory depletion)

(ii) When Y = 520,

AD = 40 + (0.8 x 520) = 40 + 416 = 456

Unplanned inventory per day = (456 - 380) / 365 = 76/365 = 0.2082 (a negative value means inventory accumulation)

(2)

Real Change in exchange rate = Nominal devaluation - Change in price level

When Real Change in exchange rate < 0, a real devaluation takes place.

(i) When Price level rises by 11%,

Real Change in exchange rate = 13% - 11% = 2%

Since Real Change in exchange rate is positive, a real devaluation has not been achieved.

(ii) When Price level rises by 16%,

Real Change in exchange rate = 13% - 16% = -3%

Since Real Change in exchange rate is negative, a real devaluation has been achieved.